# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v2

@@ 1 Introduction
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs. The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we give two applications of the main results. In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
@@ 2.1 Three types of metric mean dimensions
In this subsection, we recall the definitions of upper capacity metric mean dimension on subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension on subsets [YCZ22].
Let $(X,d,f)$ be a dynamical system. For $n\in\mathbb{N}$, we define the Bowen metric $d_{n}$ on $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\}$
for any $x,y\in X$. Then the Bowen open ball and closed ball of $x$ in the metric $d_{n}$ with radius $\epsilon$ are respectively given by
@@
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}.$
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if $x,y\in E$ with $x\neq y$ implies $d_{n}(x,y)>\epsilon$. Let $s_{n}(Z,\epsilon)$ denote the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
@@
Analogous to the definition of box dimension in fractal geometry, we can formulate the definition for upper capacity metric mean dimension of subsets using separated sets.
@@
We define the upper capacity metric mean dimension of $Z$ as
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$ (2.1)
When $Z=X$, we write $h_{top}(f,X,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)=\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively. In this case, it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00]. Such a formulation allows us to extend the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Recall that the upper capacity topological entropy of $Z$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
Especially, for any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity. Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite (upper capacity) topological entropy as $\epsilon$ tends to 00. More precisely, for sufficiently $\epsilon>0$, we may think of
@@
It is easy to see that any finite entropy system has zero metric mean dimension, and hence verifies that metric mean dimension is a good candidate to capture the dynamics of infinite entropy systems.
Similar to the definitions of Hausdorff dimension and packing dimension in fractal geometry, as an analogue of the classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which is defined by Carathéodory-Pesin structures, the corresponding dimensional characterizations of metric mean dimensions via Carathéodory-Pesin structures are Bowen and packing metric mean dimensions, which has been introduced in [LV21, YCZ22, W21] and studied by several authors in [BR23, CLS21, LL24, FKO24, YCZ24].
@@
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}$ such that $Z\subset\cup_{i}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So we set
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon).$
There is a critical value of parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
We define the Bowen upper metric mean dimension of $Z$ as
@@
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i$.
@@
There is a critical value of parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
@@
We define the packing upper metric mean dimension of $Z$ as
@@
$(1)$ If $Z_{1}\subset Z_{2}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z_{1},d)$ $\leq\overline{\rm mdim}_{M}^{UC}(f,Z_{2},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{B}_{M}(f,Z_{2},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{P}_{M}(f,Z_{2},d).$
@@
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{UC}(f,Z_{i},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{B}_{M}(f,Z_{i},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{P}_{M}(f,Z_{i},d).$
@@
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$; if $Z$ is $f$-invariant and compact, then the equalities hold.
The metric mean dimension of a given subset is in general not easy to calculate its precise value. However, the Entropy Distribution Principle allows us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems. Consequently, the powerful tool gives the lower bounds of entropy and metric mean dimension of the fractal-like sets.
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$ and a constant $C>0$ satisfying:
$(1)$ for every $n$, $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for each Bowen ball $B_{n}(x,\epsilon)$ such that $B_{n}(x,\epsilon)\cap Z\neq\emptyset$;
$(2)$ $\mu(Z)>0$,
then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
$\epsilon$-packing topological entropy also admits an analogous Entropy Distribution Principle as follows:
@@
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
$(1)$ $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$;
$(2)$ $\mu(Z)>0$,
then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
@@
For sufficiently small $\delta>0$, by definition it is enough to prove that
@@
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$. For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
$E\subset\bigcup_{j}B_{n_{i}}(x_{j},\epsilon).$
Hence by condition $(1)$, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. ∎
2.2 Measure-theoretic $\epsilon$-entropy
@@
Let $C(X)$ denote the space of continuous functions on $X$ equipped with the supermum norm. The Riesz representation theorem on compact metric space gives us an approach to endow with the weak-topology on $M(X)$, which is metrizable by the following metric
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$
where $\{\phi_{k}\}$ is a family of countable dense subset of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$. The partition entropy of $\xi$ w.r.t. $\mu$ is defined by
@@
Then we define measure-theoretic entropy of $f$ w.r.t. $\xi$ as
@@
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $f$ w.r.t. $\mu$ is defined as
@@
where the supremum is taken over all finite Borel measurable partitions $\xi$ of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions on subsets is the measure-theoretic metric mean dimensions of invariant measures. Unlike topological entropy, measure-theoretic entropy of ergodic measures admit several equivalent definitions by considering the limits of the measure-theoretic $\epsilon$-entropies: for instance, Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ24], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ24, Theorem 1.1]). Here, we only recall two types of them.
@@
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu$ is given by
@@
    The (lower) Katok $\epsilon$-entropy of $\mu$ is given by
@@
The relation between Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy is related by the following inequalities:
@@
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})<4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
@@
The next variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
@@
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is given in [YCZ24, Theorems 1.2 and 1.3].

### v3

@@ 1 Introduction
    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by making some modifications to the current proofs. The novelty of our work not only gives a proper and full extension of the previous work in [PS07, ZCC12, ZC13, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for studying the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some possible substitutions of the uniform separation property so that all the above results are true for $g$-almost product property, or even for cases beyond $\mathbb{Z}$.
The paper is organized as follows. In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In section 4, we consider some applications of the main results, and apply it to investigate the level sets of Birkhoff average of a continuous potential, Li-Yorke chaos and the metric mean dimensions of generic points of the full shift over compact metric spaces.
@@ 2.1 Three types of metric mean dimensions
In this subsection, we review the definitions of upper capacity metric mean dimension of subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension [YCZ22] of subsets.
Let $(X,d,f)$ be a t.d.s.. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on the space $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\},$
where $x,y\in X$. Then the Bowen open ball and closed ball centered at $x$ with radius $\epsilon$ in the metric $d_{n}$ are given by
@@
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
respectively.
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$. Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
@@
Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, i.e.,
Specifically, topological entropy fails to distinguish different levels of topological complexity within the infinite entropy systems
Analogous to the definition of box dimension in fractal geometry, we formulate the refined entropy-like quantity for infinite entropy systems using $\epsilon$-upper capacity topological entropy.
@@
The upper capacity metric mean dimension of $Z$ is defined by
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$
In the case where $Z$ is equal to $X$, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$ respectively. Then the upper capacity metric mean dimension of $X$ is reduced to Lindenstrauss-Weiss’s definition introduced in [LW00]. Such a formulation extends the concept of the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Notice that the (upper capacity) metric mean dimension is defined by detecting how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon$ tends to 00. That is, for sufficiently small $\epsilon>0$, we may think of
@@
It is straightforward to observe that any system with finite topological entropy has zero metric mean dimension. This, in turn, shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
The classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogue of the Hausdorff dimension and packing dimension in fractal geometry. The corresponding dimensional characterizations of metric mean dimensions, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the Bowen and packing metric mean dimensions. One may refer to [BR23, CLS21, LL24, FKO24, YCZ25] for the results involving Bowen and packing metric mean dimensions.
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i\in I}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases, this ensures that the limit
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon)$
exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)=\inf\{s>0:M(Z,s,\epsilon)=0\}=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
@@
The Bowen upper metric mean dimension of $Z$ is defined by
@@
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i\in I}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i\in I$.
@@
There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
@@
Definition 2.3.
The packing upper metric mean dimension of $Z$ is defined by
@@
$(1)$ If $Z_{1}\subset Z_{2}\subset X$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z_{1},d)\leq\overline{\rm mdim}_{M}^{S}(f,Z_{2},d).$
@@
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then for each $S\in\{UC,B,P\}$,
$\overline{\rm mdim}_{M}^{S}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{S}(f,Z_{i},d).$
@@
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$. Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions are equal.
Given a subset of phase space, it is not easy to calculate its precise Bowen and packing metric mean dimensions. However, in the multifractal analysis of dynamical systems the Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets. Consequently, the powerful tool yields the lower bounds of metric mean dimension of the fractal-like sets. $\epsilon$-Bowen and packing topological entropies of Entropy Distribution Principles are stated as follows:
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and for every $n$ and for each Bowen ball $B_{n}(x,\epsilon)$ with $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, one has $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$, then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
@@
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ that converges to $\infty$ satisfying $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
@@
By definition, it is enough to prove that for sufficiently small $\delta>0$,
@@
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$.
For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of the set $E$. Then, by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ Hence, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j\in J}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. This shows that $\mathcal{P}(Z,s-\delta,\frac{\epsilon}{6})=\infty$, and hence $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$ by letting $\delta\to 0$. ∎
2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
@@
Let $C(X)$ denote the space consisting of real-valued continuous functions defined on $X$, equipped with the supremum norm. By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$ (2.1)
where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural embedding from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the sense of the weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\cap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\cup_{i=1}^{k}C_{i})=0$. Then the partition entropy of $\xi$ w.r.t. $\mu$ is defined by
@@
We define the measurable partition of $\mu\in M_{f}(X)$ w.r.t. $\xi$ as
@@
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
@@
where the supremum is taken over all finite Borel measurable partitions of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions is the measure-theoretic metric mean dimensions of invariant measures by measuring the divergent rates of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ25, Theorem 1.1]). Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
@@
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu\in M_{f}(X)$ is given by
@@
    The lower Katok $\epsilon$-entropy of $\mu$ is given by
@@
The Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy are related by the following inequalities:
@@
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
@@
The following variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
@@
The results are also valid for lower metric mean dimension.
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is summarized in [YCZ25, Theorems 1.2 and 1.3].